# GuidNoise: Single-Pair Guided Diffusion for Generalized Noise Synthesis

### v1

@@ Introduction
Deep neural networks, particularly convolutional neural networks (CNN), have significantly advanced image denoising (DnCNN:zhang2017beyond; RIDNet:anwar2019real). Initially, image-denoising research relied on basic synthetic noise models, primarily trained on typical noise distributions such as Additive White Gaussian Noise (AWGN). These models perform well in controlled environments, but often struggle to generalize to noisy real-world images with complex noise patterns. To address this limitation, datasets consisting of real-world noisy images have been proposed, such as SIDD (SIDD:SIDD_2018_CVPR) and PolyU (PolyU:xu2018real). However, creating these datasets is costly because it requires an extensive collection of real noisy images and the corresponding acquisition of clean images paired with the target noisy image.
To reduce the costs associated with preparing noisy image data sets in real-world situations, various generative models have been proposed to synthesize noise by learning from real-world data. Specifically, Generative Adversarial Networks (GAN) (goodfellow2020generative), Normalizing Flow (NF) (papamakarios2021normalizing), and diffusion (ho2020denoising) models have shown promising results (NoiseFlow:abdelhamed2019noise; DANet_AKLD_KLD_PGap:yue2020dual; CycleISP:zamir2020cycleisp; C2N:jang2021c2n; PNGAN:cai2021learning; Flow-sRGB:kousha2022modeling; NeCA:fu2023srgb; NAFlow:kim2023srgb; wu2025realistic). The recent noise synthesis models utilize various techniques using camera metadata or a number of noisy/clean image pairs of the target scene to achieve higher noise modeling performance.
Specifically, a line of approaches (NoiseFlow:abdelhamed2019noise; Flow-sRGB:kousha2022modeling; NeCA:fu2023srgb) incorporates camera-specific information to generate noise that closely mimics the characteristics of particular imaging systems, while the others (DANet_AKLD_KLD_PGap:yue2020dual; NAFlow:kim2023srgb) use target noisy images paired with input clean images to capture the intricate patterns of noise distributions. These methods have demonstrated superior performance in generating high-fidelity noise, effectively capturing the complexities of actual camera outputs. However, while these methods perform successfully under the assumption of equivalent environments between training and testing, they often struggle to generalize across diverse conditions, including different devices, ISO settings, shutter speeds, and scene variations. Their reliance on specific metadata or a large number of paired images can limit their adaptation capability to broader datasets or scenarios where such detailed information is unavailable.
@@
The constraints of real-world noisy datasets have driven efforts towards generative noise modeling, aiming to overcome the dependency on real-world datasets by learning to synthesize realistic noise distributions. Generative Adversarial Networks (GAN), Normalizing Flows (NF), and diffusion models have emerged as prominent approaches in this area. Using synthesized noise from the generative model has shown the potential to improve the denoising networks (Finetune_w_PNGAN:cai2021learning).
DANet (DANet_AKLD_KLD_PGap:yue2020dual) simultaneously generates and removes noise, learning the joint distribution between clean and noisy image pairs from the target scene. CycleISP (CycleISP:zamir2020cycleisp) designs the bidirectional camera imaging pipeline, generating realistic image pairs for RAW and sRGB denoising. C2N (C2N:jang2021c2n) employs adversarial loss using unpaired noisy-clean images, which allows for training without relying on paired data. However, its modeling from unpaired data often suffers from limitations in noise quality, such as artifacts and color-shift problems. PNGAN (PNGAN:cai2021learning) tackles the noisy image generation task by breaking it down into image-domain and noise-domain alignment problems using pixel-level modeling. Flow-sRGB (Flow-sRGB:kousha2022modeling) extends the normalizing flow-based NoiseFlow (NoiseFlow:abdelhamed2019noise) to the sRGB space, offering improved noise modeling across different image formats. NeCA (NeCA:fu2023srgb) explicitly consists of multiple specialized noise networks based on neighbor-noise correlations. NAFlow (NAFlow:kim2023srgb) builds upon this by mapping noise from various camera models into a unified latent space.
Diffusion models (DDPM_Diffusion:ho2020denoising; DDIM_Diffusion:songdenoising; Pred-V_Diffusion:salimansprogressive) have gained significant interest for their ability to represent complex distributions without the mode collapse observed in GAN, and without the restrictive invertibility requirements of NF models. Unlike NF models, which require invertible transformations and struggle with high-dimensional complex distributions, diffusion models provide the flexibility to better capture intricate noise patterns. RNSD (wu2025realistic) is one of the first studies demonstrating the effectiveness of diffusion models in handling the complexities of real-world noise. Additionally, revisiting the dataset preparation pipeline is recently proposed (li2025noise). GuidNoise generates high-quality noisy images without requiring camera metadata or a paired noisy image corresponding to a target clean image, thereby enhancing generalizability.
@@
The core idea of diffusion model (DPM_Diffusion:sohl2015deep) involves gradually adding noise to an image through a forward diffusion process and then learning a reverse process to reconstruct the original image. Especially, we leverage Denoising Diffusion Probabilistic Models (DDPM) (DDPM_Diffusion:ho2020denoising) for diffusion modeling. In the DDPM, the forward process progressively corrupts a target image with Gaussian diffusion noise, as:
@@
To reconstruct the original image from the noisy counterpart, DDPM directly predicts the noise component added during the forward process with two subsequent improvements. Denoising Diffusion Implicit Models (DDIM) (DDIM_Diffusion:songdenoising) and v-prediction (Pred-V_Diffusion:salimansprogressive). DDIM provides a deterministic sampling,
@@
where $\alpha_{t}$ and $\sigma_{t}$ are time-dependent coefficients. $\alpha_{t}$ balances the diffusion noise and the target image and $\sigma_{t}$ balances between deterministic and stochastic sampling. The v-prediction (Pred-V_Diffusion:salimansprogressive) estimates the $\mathbf{v}_{t}$ instead of predicting the diffusion noise $\epsilon$, which improves stability even with fewer sampling steps (kingma2023understanding). This provides a more nuanced combination of the diffusion noise and the target image, allowing better control over the backward process, which defined as:
@@
Our noise generation model $\eta(\cdot)$ is based on a conditional U-Net architecture from DDPM. In training phase, $\eta(\cdot)$ takes $\mathbf{x}_{t}$ from forward process as an input to backward process, along with a clean image $\mathbf{c}\in\mathcal{S}^{tr}_{c}$ and a guidance pair $(\mathbf{x}_{\text{r}}$, $\mathbf{c}_{\text{r}})\in\mathcal{G}^{\text{tr}}$ as additional inputs for guiding the synthetic noisy image generation, as illustrated in Figure 2. We note that the guidance pairs are randomly selected from the samples in $\mathcal{G}^{tr}$ at the training phase.
@@
Inspired by the conditional feature modification methods (perez2018film; dhariwal2021diffusion; wu2025realistic), applying the affine transform of the features given the conditional information, we propose a guidance-aware affine feature modification method to embed the domain-specific noise information from the guidance pair ($\mathbf{x}_{\text{r}}$, $\mathbf{c}_{\text{r}}$). Here, the embedding $\mathbf{z_{r}}$ from guidance image pair ($\mathbf{x}_{\text{r}}$, $\mathbf{c}_{\text{r}}$) is an input into the Noise-aware guidance module $\tau$ that processes $\mathbf{z_{r}}$ to form the guidance embedding $\mathbf{h}$,
@@
Specifically, during the sampling process from T to 0, we track the gradients for the last $T_{\text{split}}$ steps, where most of the fine-grained noise details are synthesized. The refine loss aims to improve the similarity between the histogram of the synthesized noisy image and the ground truth noisy image. To this end, we use Kullback-Leibler Divergence (KLD) $D_{\text{KL}}$ and differentiable histogram (HistLoss:UstinovaNIPS16) $\bar{H}$. The refine loss is given by
@@
We compare the similarity of noisy images with several noise synthesis methods: C2N (C2N:jang2021c2n), Flow-sRGB (Flow-sRGB:kousha2022modeling), NeCA-S/NeCA-W (NeCA:fu2023srgb), and NAFlow (NAFlow:kim2023srgb). We note that C2N, Flow-sRGB, and NAFlow, as lightweight models, exhibit performance limitations, whereas NeCA includes models with varying parameter sizes, as detailed in the Supplementary Material. The parameter size of the noise synthesis module in our method is of the same order as NeCA. Also, while other synthesis methods have fewer parameters, they demonstrate inferior synthesis quality. To evaluate the quality of synthetic noisy images, we use KLD and AKLD (DANet_AKLD_KLD_PGap:yue2020dual) as metrics. Furthermore, we train DnCNN (DnCNN:zhang2017beyond) from scratch using synthetic datasets to indirectly evaluate the synthetic noisy images. We then train NAFNet (NAFNet:chen2022simple) via a self-augmentation to evaluate enhanced denoising performance. For denoising evaluation, we use PSNR and SSIM metrics.
@@
We primarily use the real-world noise dataset SIDD (SIDD:SIDD_2018_CVPR). For GuidNoise, we use the SIDD-Medium dataset split into 256 × 256 patches. To train DnCNN from scratch, we use subset of SIDD-Medium dataset split into 512 × 512 patches, following the experimental settings in  (NeCA:fu2023srgb) and  (NAFlow:kim2023srgb). We evaluate synthesized noisy images using SIDD-Validation (SIDD:SIDD_2018_CVPR). To show our model’s generalized performance, we use SIDD+ (SIDD+:abdelhamed2020ntire), PolyU (PolyU:xu2018real), and Nam (NAM:nam2016holistic) for image synthesis, and SIDD-Benchmark [^1] Our results may slightly differ from prior SIDD benchmarks due to its recent migration to a Kaggle competition. (SIDD:SIDD_2018_CVPR) for image denoising. We crop PolyU and Nam datasets to 512 × 512 pixels. For self-augmentation, we sample the training set from the SIDD-Validation using proportions of 1/16, 1/8, 1/4, and 1/2. with the remaining one as test set.
@@
GuidNoise is trained with a learning rate of 1e-5 and batch size of 4, using the diffusion loss defined in Eq. (5) for 300K iterations. The additional refining process was conducted by a batch size of 1, incorporating both the diffusion loss and the refining loss from Eq. (12) for 50K iterations. All optimization in GuidNoise is performed using AdamW (AdamW:loshchilov2018decoupled). We set $\lambda{=}0.1$, $\gamma{=}0.1$, $T_{\text{split}}{=}2$, and T=50T{=}50. To train DnCNN (DnCNN:zhang2017beyond) from scratch, we base our approach on the experimental settings in  (NeCA:fu2023srgb) and  (NAFlow:kim2023srgb). We use a learning rate of 1e-3 and run the training for 300 epochs, with a batch size of 32 and random cropping. For optimization, we employ the Adam optimizer (Adam:kingma2014adam). To train NAFNet from scratch via self-augmentation, we use a learning rate $\eta_{0}$ of 1e-3 with an exponential scheduler $(\eta_{x}=\max(0.95^{x}\cdot\eta_{0},1\times 10^{-6}))$, which decays the learning rate by 0.95 every 100 iterations initially, then every 1K iterations. We use a batch size of 16 and the PSNR Loss from NAFNet experiments (NAFNet:chen2022simple) for 10K iterations.
@@ Discussion
To see the effectiveness of the proposed noise guidance and refine loss, we demonstrate the ablation in Table 5. The guidance image reduces KLD from 0.080 to 0.050 and AKLD from 0.150 to 0.118 compared to the baseline which uses only input clean images. The refine loss further reduces KLD from 0.050 to 0.014. These results demonstrate that our cascade decoding architecture and noise-aware guidance significantly improve noise synthesis. Furthermore, a recent study PUCA (jang2023puca) trained in SIDD-Medium achieved an unsupervised performance of PSNR 37.49, slightly higher than our self-augmentation($\times 1/8$) PSNR 37.48. While PUCA uses 320 images, our approach employs a smaller subset ($\times 1/8$) from SIDD-Validation, not SIDD-Medium, so that it consists of only five pairs.
@@ Acknowledgement
Figure 1: Architecture overview of GuidNoise.
Figure 2: Architecture of Down/Up Blocks with skip connection.
Figure 3: Architecture of Residual/Attention Blocks.
Model architecture and size
The proposed reference-guided noise synthesis diffusion (GuidNoise) is based on a conditional U-Net architecture from DDPM (DDPM_Diffusion:ho2020denoising). The detailed architecture of our model is shown in Figure 1, Figure 2, and Figure 3. Specifically, the encoding process in Eq. (7) of the main manuscript corresponds to the Encoder in Figure 1. The guidance-aware affine feature modification (GAFM) $\tau$ in Eq. (8) corresponds to the Conditioner in Figure 2. The cascade decoding process in Eq. (9) corresponds to the decoding process in Figure 3. The number of parameters for GuidNoise and others can be compared as shown in Table 2, where we note that lightweight models such as C2N, Flow-sRGB, and NAFlow have performance limitations, while NeCA includes models with varying parameter sizes.
Self-augmentation
Model Variants GuidNoise allows us to augment limited real noise datasets, which is referred to as Self-augmentation. We conduct the experiments for self-augmented training across various dataset size and model size of NAFNet (NAFNet:chen2022simple). The specific model configuration of each size in Table 1.
Results Due to the varying results depending on model size and data size, we select representative trend results as shown in Figure 5 in the main manuscript. We present all results, including the remaining trends on Figure 4. Self-augmentation via GuidNoise shows significant benefits, especially with limited training data and smaller model architectures.
Model Type Channel Width #Encoder Blocks #Middle Blocks #Decoder Blocks
Tiny 8 1, 1, 1, 2 3 1, 1, 1, 1
Small 12 1, 1, 2, 3 4 1, 1, 1, 1
Medium 32 2, 2, 4, 8 12 2, 2, 2, 2
Large 64 2, 2, 4, 8 12 2, 2, 2, 2
Table 1: Configurations of NAFNet variants.
Model C2N Flow-sRGB NAFlow
#Params 2.2M 6K 1.1M
Model NeCA-S NeCA-W Ours
#Params $5\times(7.8\text{M}^{c})$ $5\times(7.8\text{M}^{c}+\alpha)$ 62M
Table 2: Comparison of model parameters. The remaining parameter size of NeCA is set to $\alpha=263\text{K}^{a}+42\text{K}^{b}$.
(a) PSNR comparison across dataset sizes on NAFNet-Tiny
(b) PSNR comparison across model sizes on $\times 1/16$ data
(c) PSNR comparison across dataset sizes on NAFNet-Small
(d) PSNR comparison across model sizes on $\times 1/8$ data
(e) PSNR comparison across dataset sizes on NAFNet-Medium
(f) PSNR comparison across model sizes on $\times 1/4$ data
(g) PSNR comparison across dataset sizes on NAFNet-Large
(h) PSNR comparison across model sizes on $\times 1/2$ data
Figure 4: Denoising performance trends across model size (left) and dataset size (right). The smaller the training data or the smaller the model, the larger the self-augmentation effect tends to be.
Dataset Size NAFNet-Tiny NAFNet-Small NAFNet-Medium NAFNet-Large
Real dataset Self-augmented Real dataset Self-augmented Real dataset Self-augmented Real dataset Self-augmented
$\uparrow PSNR\uparrow$ $\uparrow PSNR\uparrow$ $\uparrow PSNR\uparrow$ $\uparrow PSNR\uparrow$
$\times$1/16 33.80 35.12 34.19 35.83 35.48 36.42 36.54 36.91
$\times$1/8 34.82 36.22 34.84 36.62 36.10 37.06 36.96 37.48
$\times$1/4 35.61 36.93 35.79 37.21 36.60 37.48 37.46 37.82
$\times$1/2 36.54 37.43 36.65 37.61 37.08 37.76 37.50 38.01
#Parameters 0.56M 1.59M 29.16M 115.98M
Table 3: Denoising performance comparison across model and dataset sizes. PSNR values on SIDD-Validation for different model configurations (Tiny, Small, Medium, Large) and dataset sizes (×1/16, ×1/8, ×1/4, ×1/2) trained on real dataset and self-augmented datasets.
(a) wo/ Refine Loss
(b) w/ Refine Loss
(c) Noise distribution comparison
Figure 6: Noise Distribution Visualization. Figures 6(a) and  6(b) show the noise distribution changes per time step. Figure 6(c) contrasts noise distributions of real noisy data, images generated with refine loss (w/ Refine), and without (wo/ Refine).
Noisy
C2N
NeCA-W
NAFlow
Ours
Clean
Figure 7: Qualitative comparison of denoising results on the SIDD-Validation dataset.
Refine Loss
We validate the proposed refine loss that refines the diffusion process to consider the noise distribution of synthesized images. As shown in Figure 6, the synthesized noise distribution at Step 0 with the refine loss looks similar to the real noise distribution, while that without the refine loss can be seen as very different from the real noise distribution. During DDIM sampling process, the noise distribution primarily undergoes large-scale adjustments in most time steps, mainly in the form of shifts and scaling. However, fine adjustments occur within the final 1–2 steps. Leveraging this insight, we define the refine loss and set the differentiable sampling step $T_{\text{split}}$ to last two sampling steps.
Qualitative denoising results
We present the visualization of denoised noisy images in Figure 7. Notably, the denoised images from the denoising model trained with synthesized images from GuidNoise are visually closest to the clean image. This indirectly supports the realistic synthesis performance of GuidNoise.

### v2

@@ Introduction
Deep neural networks, particularly convolutional neural networks (CNN), have significantly advanced image denoising (Zhang et al. 2017; Anwar and Barnes 2019). Initially, image-denoising research relied on basic synthetic noise models, primarily trained on typical noise distributions such as Additive White Gaussian Noise (AWGN). These models perform well in controlled environments, but often struggle to generalize to noisy real-world images with complex noise patterns. To address this limitation, datasets consisting of real-world noisy images have been proposed, such as SIDD (Abdelhamed et al. 2018) and PolyU (Xu et al. 2018). However, creating these datasets is costly because it requires an extensive collection of real noisy images and the corresponding acquisition of clean images paired with the target noisy image.
To reduce the costs associated with preparing noisy image data sets in real-world situations, various generative models have been proposed to synthesize noise by learning from real-world data. Specifically, Generative Adversarial Networks (GAN) (Goodfellow et al. 2020), Normalizing Flow (NF) (Papamakarios et al. 2021), and diffusion (Ho et al. 2020b) models have shown promising results (Abdelhamed et al. 2019; Yue et al. 2020; Zamir et al. 2020; Jang et al. 2021; Cai et al. 2021b; Kousha et al. 2022; Fu et al. 2023; Kim et al. 2024; Wu et al. 2025). The recent noise synthesis models utilize various techniques using camera metadata or a number of noisy/clean image pairs of the target scene to achieve higher noise modeling performance.
Specifically, a line of approaches (Abdelhamed et al. 2019; Kousha et al. 2022; Fu et al. 2023) incorporates camera-specific information to generate noise that closely mimics the characteristics of particular imaging systems, while the others (Yue et al. 2020; Kim et al. 2024) use target noisy images paired with input clean images to capture the intricate patterns of noise distributions. These methods have demonstrated superior performance in generating high-fidelity noise, effectively capturing the complexities of actual camera outputs. However, while these methods perform successfully under the assumption of equivalent environments between training and testing, they often struggle to generalize across diverse conditions, including different devices, ISO settings, shutter speeds, and scene variations. Their reliance on specific metadata or a large number of paired images can limit their adaptation capability to broader datasets or scenarios where such detailed information is unavailable.
@@
The constraints of real-world noisy datasets have driven efforts towards generative noise modeling, aiming to overcome the dependency on real-world datasets by learning to synthesize realistic noise distributions. Generative Adversarial Networks (GAN), Normalizing Flows (NF), and diffusion models have emerged as prominent approaches in this area. Using synthesized noise from the generative model has shown the potential to improve the denoising networks (Cai et al. 2021a).
DANet (Yue et al. 2020) simultaneously generates and removes noise, learning the joint distribution between clean and noisy image pairs from the target scene. CycleISP (Zamir et al. 2020) designs the bidirectional camera imaging pipeline, generating realistic image pairs for RAW and sRGB denoising. C2N (Jang et al. 2021) employs adversarial loss using unpaired noisy-clean images, which allows for training without relying on paired data. However, its modeling from unpaired data often suffers from limitations in noise quality, such as artifacts and color-shift problems. PNGAN (Cai et al. 2021b) tackles the noisy image generation task by breaking it down into image-domain and noise-domain alignment problems using pixel-level modeling. Flow-sRGB (Kousha et al. 2022) extends the normalizing flow-based NoiseFlow (Abdelhamed et al. 2019) to the sRGB space, offering improved noise modeling across different image formats. NeCA (Fu et al. 2023) explicitly consists of multiple specialized noise networks based on neighbor-noise correlations. NAFlow (Kim et al. 2024) builds upon this by mapping noise from various camera models into a unified latent space.
Diffusion models (Ho et al. 2020a; Song et al. 2021; Salimans and Ho 2022) have gained significant interest for their ability to represent complex distributions without the mode collapse observed in GAN, and without the restrictive invertibility requirements of NF models. Unlike NF models, which require invertible transformations and struggle with high-dimensional complex distributions, diffusion models provide the flexibility to better capture intricate noise patterns. RNSD (Wu et al. 2025) is one of the first studies demonstrating the effectiveness of diffusion models in handling the complexities of real-world noise. Additionally, revisiting the dataset preparation pipeline is recently proposed (Li et al. 2025). GuidNoise generates high-quality noisy images without requiring camera metadata or a paired noisy image corresponding to a target clean image, thereby enhancing generalizability.
@@
The core idea of diffusion model (Sohl-Dickstein et al. 2015) involves gradually adding noise to an image through a forward diffusion process and then learning a reverse process to reconstruct the original image. Especially, we leverage Denoising Diffusion Probabilistic Models (DDPM) (Ho et al. 2020a) for diffusion modeling. In the DDPM, the forward process progressively corrupts a target image with Gaussian diffusion noise, as:
@@
To reconstruct the original image from the noisy counterpart, DDPM directly predicts the noise component added during the forward process with two subsequent improvements. Denoising Diffusion Implicit Models (DDIM) (Song et al. 2021) and v-prediction (Salimans and Ho 2022). DDIM provides a deterministic sampling,
@@
where $\alpha_{t}$ and $\sigma_{t}$ are time-dependent coefficients. $\alpha_{t}$ balances the diffusion noise and the target image and $\sigma_{t}$ balances between deterministic and stochastic sampling. The v-prediction (Salimans and Ho 2022) estimates the $\mathbf{v}_{t}$ instead of predicting the diffusion noise $\epsilon$, which improves stability even with fewer sampling steps (Kingma and Gao 2023). This provides a more nuanced combination of the diffusion noise and the target image, allowing better control over the backward process, which defined as:
@@
Our noise generation model $\eta(\cdot)$ is based on a conditional U-Net architecture from DDPM. In training phase, $\eta(\cdot)$ takes $\mathbf{x}_{t}$ from forward process as an input to backward process, along with a clean image $\mathbf{c}\in\mathcal{S}^{tr}_{c}$ and a guidance pair $(\mathbf{x}_{\text{r}}$, $\mathbf{c}_{\text{r}})\in\mathcal{G}^{\text{tr}}$ as additional inputs for guiding the synthetic noisy image generation, as illustrated in Figure 2.
@@
Inspired by the conditional feature modification methods (Perez et al. 2018; Dhariwal and Nichol 2021; Wu et al. 2025), applying the affine transform of the features given the conditional information, we propose a guidance-aware affine feature modification method to embed the domain-specific noise information from the guidance pair ($\mathbf{x}_{\text{r}}$, $\mathbf{c}_{\text{r}}$). Here, the embedding $\mathbf{z_{r}}$ from guidance image pair ($\mathbf{x}_{\text{r}}$, $\mathbf{c}_{\text{r}}$) is an input into the Noise-aware guidance module $\tau$ that processes $\mathbf{z_{r}}$ to form the guidance embedding $\mathbf{h}$,
@@
Specifically, during the sampling process from T to 0, we track the gradients for the last $T_{\text{split}}$ steps, where most of the fine-grained noise details are synthesized. The refine loss aims to improve the similarity between the histogram of the synthesized noisy image and the ground truth noisy image. To this end, we use Kullback-Leibler Divergence (KLD) $D_{\text{KL}}$ and differentiable histogram (Ustinova and Lempitsky 2016) $\bar{H}$. The refine loss is given by
@@
We compare the similarity of noisy images with several noise synthesis methods: C2N (Jang et al. 2021), Flow-sRGB (Kousha et al. 2022), NeCA-S/NeCA-W (Fu et al. 2023), and NAFlow (Kim et al. 2024). We note that C2N, Flow-sRGB, and NAFlow, as lightweight models, exhibit performance limitations, whereas NeCA includes models with varying parameter sizes, as detailed in the Supplementary Material. The parameter size of the noise synthesis module in our method is of the same order as NeCA. Also, while other synthesis methods have fewer parameters, they demonstrate inferior synthesis quality. To evaluate the quality of synthetic noisy images, we use KLD and AKLD (Yue et al. 2020) as metrics. Furthermore, we train DnCNN (Zhang et al. 2017) from scratch using synthetic datasets to indirectly evaluate the synthetic noisy images. We then train NAFNet (Chen et al. 2022) via a self-augmentation to evaluate enhanced denoising performance. For denoising evaluation, we use PSNR and SSIM metrics.
@@
We primarily use the real-world noise dataset SIDD (Abdelhamed et al. 2018). For GuidNoise, we use the SIDD-Medium dataset split into 256 × 256 patches. To train DnCNN from scratch, we use subset of SIDD-Medium dataset split into 512 × 512 patches, following the experimental settings in  (Fu et al. 2023) and  (Kim et al. 2024). We evaluate synthesized noisy images using SIDD-Validation (Abdelhamed et al. 2018). To show our model’s generalized performance, we use SIDD+ (Abdelhamed et al. 2020), PolyU (Xu et al. 2018), and Nam (Nam et al. 2016) for image synthesis, and SIDD-Benchmark [^1] Our results may slightly differ from prior SIDD benchmarks due to its recent migration to a Kaggle competition. (Abdelhamed et al. 2018) for image denoising. We crop PolyU and Nam datasets to 512 × 512 pixels. For self-augmentation, we sample the training set from the SIDD-Validation using proportions of 1/16, 1/8, 1/4, and 1/2. with the remaining one as test set.
@@
GuidNoise is trained for 300K iterations using the diffusion loss in Eq. (5), with a learning rate of $1\mathrm{e}{-4}$ and a batch size of 4. It is then refined for 50K iterations with a learning rate of $1\mathrm{e}{-5}$ and a batch size of 1, using both the diffusion loss and the refining loss as defined in Eq. (12). All optimization in GuidNoise is performed using AdamW (Loshchilov and Hutter 2019). We set $\lambda{=}1$, $\gamma{=}0.1$, $T_{\text{split}}{=}2$, and T=50T{=}50. To train DnCNN (Zhang et al. 2017) from scratch, we base our approach on the experimental settings in  (Fu et al. 2023) and  (Kim et al. 2024). We use a learning rate of $1\mathrm{e}{-3}$ and run the training for 300 epochs, with a batch size of 32 and random cropping. For optimization, we employ the Adam optimizer (Kingma 2014). To train NAFNet from scratch via self-augmentation, we use a learning rate $\eta_{0}$ of $1\mathrm{e}{-3}$ with an exponential scheduler $(\eta_{x}=\max(0.95^{x}\cdot\eta_{0},1\times 10^{-6}))$, which decays the learning rate by 0.95 every 100 iterations initially, then every 1K iterations. We use a batch size of 16 and the PSNR Loss from NAFNet experiments (Chen et al. 2022) for 10K iterations.
@@ Discussion
To see the effectiveness of the proposed noise guidance and refine loss, we demonstrate the ablation in Table 5. The guidance image reduces KLD from 0.080 to 0.050 and AKLD from 0.150 to 0.118 compared to the baseline which uses only input clean images. The refine loss further reduces KLD from 0.050 to 0.014. These results demonstrate that our cascade decoding architecture and noise-aware guidance significantly improve noise synthesis. Furthermore, a recent study PUCA (Jang et al. 2023) trained in SIDD-Medium achieved an unsupervised performance of PSNR 37.49, slightly higher than our self-augmentation($\times 1/8$) PSNR 37.48. While PUCA uses 320 images, our approach employs a smaller subset ($\times 1/8$) from SIDD-Validation, not SIDD-Medium, so that it consists of only five pairs.